# TE-PAI: Exact Time Evolution by Sampling Random Circuits

### v2

@@ IV.1 Method 1: direct gate synthesis
The most straightforward approach is a direct gate synthesis whereby the non-Clifford rotation is decomposed into a sequence of Clifford gates and typically T-gates. Here we consider approximating $R_{Z}(\Delta)$ to a precision $\epsilon$ deterministically without using ancilla qubits as in [46], which requires $\approx 3.02\log_{2}(1/\epsilon)+1.77$ T-gates on average. Thus, in our example we estimate approximately $62$ T-gates are required to synthesize each rotation gate to a precision $\epsilon=10^{-6}$ which adds up to a total of $2,438,336$ T-gates. In contrast, performing Trotterisation with $N=10000$ rounds requires $4,000,000$ rotation gates each with a precision $\epsilon=10^{-8}$ which adds up to a total of $356,000,000$ T-gates. As a remark, by using approximate or non-deterministic synthesis, e.g., repeat-until-success synthesis, the costs can be reduced to $\approx 1.03\log_{2}(1/\epsilon)+5.75$ [47, 48, 49]. In contrast, exact synthesis can also be achieved by randomly choosing from a library of short T-depth approximate rotations [39].
@@ IV.2 Repeat-Until-Success Methods
|$\ket{\theta}\equiv R_{Z}(\theta)|+\rangle=\frac{1}{\sqrt{2}}\left(e^{-i\theta/2}|0\rangle+e^{+i\theta/2}|1\rangle\right).$
With probability $1/2$ the teleportation circuit yields a measurement outcome $+1$ which indicates that the qubit is correctly rotated as $R_{Z}(\theta)|\psi\rangle$. However, with equal chance, it yields a $-1$ measurement outcome which indicates an inverse rotation $R_{Z}(-\theta)|\psi\rangle$. In the latter case, one needs to apply a rotation gate with twice the angle $R_{Z}(2\theta)$ in order to obtain the desired $R_{Z}(\theta)|\psi\rangle$ net effect. The approach is therefore repeated iteratively using resource states with angle settings $2^{k}\theta$ until a $+1$ outcome is achieved, which in general requires on average the following number of trials as
$21\times\frac{1}{2}+2\times\frac{1}{4}+3\times\frac{1}{8}+\cdots=\sum_{i=1}^{\infty}\frac{i}{2^{i}}=$ 2 .
Using the Clifford hierarchy, which we define in Appendix F, is particularly beneficial for our purposes. In particular, in the following we assume that our rotation angles are of the form $\Delta_{\ell}=\pi 2^{-\ell+1}$, thus the rotation gate $R_{Z}(\Delta_{\ell})$ is in the $\ell$-th Clifford hierarchy. The significant advantage is that the repeat until success approach can terminate successfully after $\ell{-}3$ unsuccessful trials given the $\ell=3$ case is a T-gate which can be applied deterministically using T-state teleportation. Since we consider $\Delta\equiv\Delta_{\ell_{0}=9}$ in the present example, the probability of termination with a T state teleportation is relatively high as $2^{-6}$.
We note that we can directly prepare Clifford hierarchy states |$\ket{\Delta_{\ell}}$ by distillation using Reed-Mueller codes [50]. This may be beneficial when $\ell$ is sufficiently low, however for large $\ell$, the cost of directly distilling a resource state |$\ket{\Delta_{\ell}}$ to high precision will likely exceed the cost of distilling a multiple T-states that can produce |$\ket{\Delta_{\ell}}$ using the techniques we now introduce. Therefore, we consider alternative approaches that enable the preparation of these resource states by consuming relatively few T-states [50].
@@ IV.2.1 Method 2: Hamming weight phasing
Using the Hamming weight phasing method introduced in [51] and Appendix A of [52], we can efficiently apply $n$ equal-angle $R_{z}(\theta)$ rotations simultaneously to $n$ qubits. The technique still uses $\lfloor\log_{2}n+1\rfloor$ directly synthesised, arbitrary-angle $R_{z}(\cdot)$ rotations but only uses in addition at most $4(n-1)$ T-gates (or $n-1$ Toffoli gates) and $n-1$ ancilla qubits. The total required number of T-gates is therefore
@@
While we could directly apply these rotations to the computational qubits, instead we apply them to $n$ storage qubits to prepare resource states for the repeat-until-success approach – then the more storage qubits are available, the more T-gate savings this approach can yield and in the limit of infinite storage space the approach could produce $K$ resource states using $4(K-1)$ T-states.
We assume that in the present example we store $n_{\ell}=2^{\ell-4}$ number of |$\ket{\Delta_{\ell}}$ resource states, i.e., we use $n_{9}=32$ qubits to store the states |$\ket{\Delta_{\ell=9}}$ and store only one of the |$\ket{\Delta_{\ell=4}}$. We do not assume additional storage qubits for the $\ell=3$ case given this is a T-state which we assume is provided natively on demand in the fault-tolerant machine. The total storage space then adds up to $n_{9}+n_{8}+\dots n_{4}=\sum_{\ell=4}^{9}2^{\ell-4}=63$ qubits, whereas the total number of ancillary qubits is $(n_{9}-1)+(n_{8}-1)+\dots+(n_{4}-1)=57$.
@@ IV.2.2 Method 3: Catalyst generation of resource states
A catalyst tower construction was proposed in ref. [53], which builds on refs. [52, 54]. The central object is a so-called catalyst circuit that we define in Fig. 12 and which consumes two $\ket{+}$ states, a resource state |$\ket{\Delta}$ and a rotation gate $R_{z}(2\Delta)$, and outputs three resource states |$\ket{\Delta}$, thus, in effect applies two $R_{z}(\Delta)$ rotations at the cost of consuming one $R_{z}(2\Delta)$ rotation and 4 T-states. Ref. [53] then stacked these catalyst circuits so that the overall circuit prepares a family of resource states |$\ket{2^{k}\Delta}$ by catalysis consuming only a single rotation $R_{z}(2^{h}\Delta)$ where $h$ is the height of the tower.
Using our Clifford hierarchy construction with angles $\Delta\equiv\Delta_{\ell_{0}}$, the tower height can be set to $h=\ell_{0}-3$, given the rotation angle $2^{h}\Delta_{\ell_{0}}=\Delta_{3}$ can be applied directly by T-state teleportation. Thus, the catalyst towers can be initialised by the family of resource states |$\ket{\Delta_{\ell=4\dots\ell_{0}}}$ (negligible initial cost) and can then continuously produce the required resource states. Thus, we prepare and store in total $n_{\ell}=2^{\ell-4}$ resource states as in the previous subsection.
In Section F.3, we explicitly construct catalyst towers that branch out a relatively large number of catalysts at the top level to produce $n_{\ell}$ resource states |$\ket{\Delta_{\ell_{0}}}$ and branch out fewer and fewer at the lower levels, in order to precisely output the desired exponential distribution $n_{\ell}$ of the resource states |$\ket{\Delta_{\ell}}$ required for the repeat-until-success method. Then the total number of ancilla qubits required to produce the desired distribution $n_{\ell}=2^{\ell-4}$ is
@@ V Conclusion and Discussion
We introduced TE-PAI to estimate observable expected values from effectively exactly time-evolved quantum states. The approach proceeds by generating a number of random circuits in classical pre-processing, whose outputs are post-processed to yield on average exact time evolution. A significant advantage of the approach is that the random circuits are built entirely of Pauli rotations $R_{\sigma}(\cdot)$ using the Pauli operators $\sigma$ in the system Hamiltonian and using only two kinds of rotation angles $\Delta$ and $\pi$, which is particularly well-suited for fault-tolerant implementations. Furthermore, another significant advantage of TE-PAI is that it allows for a highly configurable trade-off between circuit depth and measurement overhead by adjusting a single parameter, $\Delta$, offering flexibility to fine-tune. This feature is particularly useful in NISQ and early-FTQC devices, where circuit depth and qubit coherence are the primary limitations. Furthermore, we proved that our circuits saturate the Lieb-Robinson bound in the sense that the number of gates required for simulating a total time $T$ is directly proportional to $T$.
@@
Compared to other time-evolution algorithms, the key advantage of this approach is its ability to simulate time evolution without discretisation or algorithmic errors in the sense that finite Trotterisation error can be suppressed efficiently in classical pre-processing to arbitrarily low levels without affecting circuit depth. This is a particularly powerful feature when the aim is to simulate the evolution under time-dependent Hamiltonians. Furthermore, we require no ancillary qubits or advanced quantum resources, only the ability to execute random circuits with rotation angles $\Delta$ and $\pi$, and perform measurements. The main limitation of the approach is that its measurement overhead potentially grows exponentially unless the circuit depth is increased with growing system size. Nonetheless, an approximate version of TE-PAI can interpolate between the edge cases of exact time evolution with measurement overhead and approximate time evolution but no measurement overhead by tuning a continuous trade-off parameter $\lambda$ from [39].
While TE-PAI is well-suited for NISQ applications due to its shallow depth, we develop particularly efficient fault-tolerant implementations building on the observation that the only non-Clifford gates we require are single-qubit $R_{z}(\Delta)$ rotations with identical rotation angles. Our architecture prepares resource states $R_{z}(\Delta)\ket{+}$ using less than 4 T-states via catalyst towers [53, 51] and applies the desired rotations via repeat until success teleportation [48].
@@
TE-PAI can also naturally be combined with other randomised quantum protocols. First, our random circuits could be distributed via circuit cutting algorithms and phase estimation algorithms. For instance, Ref [55] employs circuit-cutting techniques using the quasiprobabilistic decomposition framework of our method to remove entangling gates rather than to reduce overall circuit depth. Moreover,  [56] proposes a randomized algorithm for phase estimation that balances gate complexity and shot noise, aiming to eliminate dependence on the sparsity of the Hamiltonian. Second, TE-PAI can be combined immediately with classical shadows by appending a layer of random measurement-basis transformation gates [16, 17]. This opens up powerful applications, such as shadow spectroscopy [20] or subspace expansion using time-evolved trial states [57]. Alternatively, one can also combine the approach with Pauli grouping techniques [18, 19]. Third, the approach can also be combined with algorithms whereby random initial states are time evolved, such as when estimating the density of states [24]. As many of these randomised protocols treat the evolution time $T$ a random variable, TE-PAI can be used naturally to implement queries to random evolution times. Furthermore, our random circuits are composed of Pauli operators that appear in the Hamiltonian and can thus present opportunities for further optimisation through advanced compilation and transpilation tools, such as 2QAN, which was specifically designed for structurally similar Trotter circuits [34] and can significantly reduce the circuit depth by parallel execution of non-overlapping gates.
@@
Given its simplicity, our approach is immediately deployable to a broad range of problems with applications in, e.g., quantum chemistry, materials science, combinatorial optimisation, high-energy physics, etc., given time evolution is one of the most important quantum algorithmic subroutines.
@@ Acknowledgements
The authors thank Hayata Morisaki, Simon Benjamin, Gergory Boyd and Zhu Sun for helpful technical conversations. This work was supported by JST ASPIRE Japan Grant Number JPMJAP2319. BK thanks the University of Oxford for a Glasstone Research Fellowship and Lady Margaret Hall, Oxford for a Research Fellowship. The numerical modelling involved in this study made use of the Quantum Exact Simulation Toolkit (QuEST) [58] via the QuESTlink [59] frontend. BK thanks UKRI for the Future Leaders Fellowship Theory to Enable Practical Quantum Advantage (MR/Y015843/1). The authors also acknowledge funding from the EPSRC projects Robust and Reliable Quantum Computing (RoaRQ, EP/W032635/1) and Software Enabling Early Quantum Advantage (SEEQA, EP/Y004655/1).
@@ Appendix A Comparison to Prior Work
We compare TE-PAI to a related algorithm developed in ref. [29] which, similar to TE-PAI, simulates the time evolution by averaging over random quantum circuits. The approach decomposes the small-angle unitary rotations into linear combination of unitary matrices which are then randomly sampled to yield an unbiased estimator. In contrast, TE-PAI obtains an unbiased estimator by randomising a linear combination of superoperators in Eq. 8.

### v3

@@ IV.1 Method 1: direct gate synthesis
The most straightforward approach is a direct gate synthesis whereby the non-Clifford rotation is decomposed into a sequence of Clifford gates and typically T-gates. Here we consider approximating $R_{Z}(\Delta)$ to a precision $\epsilon$ deterministically without using ancilla qubits as in [synthesis0], which requires $\approx 3.02\log_{2}(1/\epsilon)+1.77$ T-gates on average. Thus, in our example we estimate approximately $62$ T-gates are required to synthesize each rotation gate to a precision $\epsilon=10^{-6}$ which adds up to a total of $2,438,336$ T-gates. In contrast, performing Trotterisation with $N=10000$ rounds requires $4,000,000$ rotation gates each with a precision $\epsilon=10^{-8}$ which adds up to a total of $356,000,000$ T-gates. As a remark, by using approximate or non-deterministic synthesis, e.g., repeat-until-success synthesis, the costs can be reduced to $\approx 1.03\log_{2}(1/\epsilon)+5.75$ [synthesis2, synthesis, synthesis_summary]. In contrast, exact synthesis can also be achieved by randomly choosing from a library of short T-depth approximate rotations [koczor2024sparse].
@@ IV.2 Repeat-Until-Success Methods
[…]
With probability $1/2$ the teleportation circuit yields a measurement outcome +1+1 which indicates that the qubit is correctly rotated as $R_{Z}(\theta)|\psi\rangle$. However, with equal chance, it yields a 1-1 measurement outcome which indicates an inverse rotation $R_{Z}(-\theta)|\psi\rangle$. In the latter case, one needs to apply a rotation gate with twice the angle $R_{Z}(2\theta)$ in order to obtain the desired $R_{Z}(\theta)|\psi\rangle$ net effect. The approach is therefore repeated iteratively using resource states with angle settings $2^{k}\theta$ until a +1+1 outcome is achieved, which in general requires on average the following number of trials as
$1\times\frac{1}{2}+2\times\frac{1}{4}+3\times\frac{1}{8}+\cdots=\sum_{i=1}^{\infty}\frac{i}{2^{i}}=2.$
Using the Clifford hierarchy, which we define in LABEL:app:resource, is particularly beneficial for our purposes. In particular, in the following we assume that our rotation angles are of the form $\Delta_{\ell}=\pi 2^{-\ell+1}$, thus the rotation gate $R_{Z}(\Delta_{\ell})$ is in the $\ell$-th Clifford hierarchy. The significant advantage is that the repeat until success approach can terminate successfully after $\ell{-}3$ unsuccessful trials given the $\ell=3$ case is a T-gate which can be applied deterministically using T-state teleportation. Since we consider $\Delta\equiv\Delta_{\ell_{0}=9}$ in the present example, the probability of termination with a T state teleportation is relatively high as $2^{-6}$.
[…]
@@ IV.2.1 Method 2: Hamming weight phasing
Using the Hamming weight phasing method introduced in [hamming-Gidney] and Appendix A of [hamming], we can efficiently apply $n$ equal-angle $R_{z}(\theta)$ rotations simultaneously to $n$ qubits. The technique still uses $\lfloor\log_{2}n+1\rfloor$ directly synthesised, arbitrary-angle $R_{z}(\cdot)$ rotations but only uses in addition at most 4(n1)4(n-1) T-gates (or n1n-1 Toffoli gates) and n1n-1 ancilla qubits. The total required number of T-gates is therefore
@@
While we could directly apply these rotations to the computational qubits, instead we apply them to $n$ storage qubits to prepare resource states for the repeat-until-success approach – then the more storage qubits are available, the more T-gate savings this approach can yield and in the limit of infinite storage space the approach could produce $K$ resource states using 4(K1)4(K-1) T-states.
[…]
@@ IV.2.2 Method 3: Catalyst generation of resource states
[…]
@@ V Conclusion and Discussion
We introduced TE-PAI to estimate observable expected values from effectively exactly time-evolved quantum states. The approach proceeds by generating a number of random circuits in classical pre-processing, whose outputs are post-processed to yield on average exact time evolution. A significant advantage of the approach is that the random circuits are built entirely of Pauli rotations $R_{\sigma}(\cdot)$ using the Pauli operators $\sigma$ in the system Hamiltonian and using only two kinds of rotation angles $\Delta$ and $\pi$, which is particularly well-suited for fault-tolerant implementations. Furthermore, another significant advantage of TE-PAI is that it allows for a highly configurable trade-off between circuit depth and measurement overhead by adjusting a single parameter, $\Delta$, offering flexibility to fine-tune. This feature is particularly useful in NISQ and early-FTQC devices, where circuit depth and qubit coherence are the primary limitations. In the limit $\Delta\to\pi/2$, TE-PAI reduces to a Clifford circuit that is classically simulable—and Clifford circuits can be trivially exectued in fault tolerant quantum computers—but requires infinite sampling overhead. In contrast, in the limit $\Delta\to 0$ the measurement overhead vanishes at the cost of an infinitely deep quantum circuit. TE-PAI thus provides a continuous interpolation between these extremes and trades off coherent resources for incorherent repetition, i.e., by tuning $\Delta$, one effectively trades classical resources (sampling) against quantum resources (circuit depth) as also relevant in quantum resource theory. Furthermore, we proved that our circuits saturate the Lieb-Robinson bound in the sense that the number of gates required for simulating a total time $T$ is directly proportional to $T$. By construction, TE-PAI respects the Lieb-Robinson bound on information propagation in expectation, so that quantities such as entanglement growth and state-transfer fidelity are reproduced on average as in the ideal evolution. These observations indicate the potential for TE-PAI’s randomized structure to be compatible with error-resilient strategies in quantum communication and computation.
@@
Compared to other time-evolution algorithms, the key advantage of this approach is its ability to simulate time evolution without discretisation or algorithmic errors in the sense that finite Trotterisation error can be suppressed efficiently in classical pre-processing to arbitrarily low levels without affecting circuit depth. This is a particularly powerful feature when the aim is to simulate the evolution under time-dependent Hamiltonians. Furthermore, we require no ancillary qubits or advanced quantum resources, only the ability to execute random circuits with rotation angles $\Delta$ and $\pi$, and perform measurements. The main limitation of the approach is that its measurement overhead potentially grows exponentially unless the circuit depth is increased with growing system size. Nonetheless, an approximate version of TE-PAI can interpolate between the edge cases of exact time evolution with measurement overhead and approximate time evolution but no measurement overhead by tuning a continuous trade-off parameter $\lambda$ from [koczor2024sparse]. Building on this flexibility, future work could explore variance reduction techniques to further reduce sampling cost.
[…]
@@
TE-PAI can also naturally be combined with other randomised quantum protocols. First, our random circuits could be distributed via circuit cutting algorithms and phase estimation algorithms. For instance, Ref [circuit_cutting] employs circuit-cutting techniques using the quasiprobabilistic decomposition framework of our method to remove entangling gates rather than to reduce overall circuit depth. Moreover,  [statistical-phase-estimation] proposes a randomized algorithm for phase estimation that balances gate complexity and shot noise, aiming to eliminate dependence on the sparsity of the Hamiltonian. Second, TE-PAI can be combined immediately with classical shadows by appending a layer of random measurement-basis transformation gates [mitigation_shadow, shadow]. This opens up powerful applications, such as shadow spectroscopy [chan2022algorithmic] or subspace expansion using time-evolved trial states [boyd2024high]. Alternatively, one can also combine the approach with Pauli grouping techniques [Crawford2019, jena2019pauli]. Third, the approach can also be combined with algorithms whereby random initial states are time evolved, such as when estimating the density of states [goh2024direct]. As many of these randomised protocols treat the evolution time $T$ a random variable, TE-PAI can be used naturally to implement queries to random evolution times. Furthermore, our random circuits are composed of Pauli operators that appear in the Hamiltonian and can thus present opportunities for further optimisation through advanced compilation and transpilation tools, such as 2QAN, which was specifically designed for structurally similar Trotter circuits [2qan] and can significantly reduce the circuit depth by parallel execution of non-overlapping gates.
@@
Given its simplicity, our approach is immediately deployable to a broad range of problems with applications in, e.g., quantum chemistry, materials science, combinatorial optimisation, high-energy physics, optimisation, quantum machine learning etc., given time evolution is one of the most important quantum algorithmic subroutines.
@@ Acknowledgements
The authors thank Hayata Morisaki, Simon Benjamin, Gergory Boyd and Zhu Sun for helpful technical conversations. This work was supported by JST ASPIRE Japan Grant Number JPMJAP2319. BK thanks the University of Oxford for a Glasstone Research Fellowship and Lady Margaret Hall, Oxford for a Research Fellowship. The numerical modelling involved in this study made use of the Quantum Exact Simulation Toolkit (QuEST) [quest] via the QuESTlink [questlink] frontend. BK thanks UKRI for the Future Leaders Fellowship Theory to Enable Practical Quantum Advantage (MR/Y015843/1). The authors also acknowledge funding from the EPSRC projects Robust and Reliable Quantum Computing (RoaRQ, EP/W032635/1) and Software Enabling Early Quantum Advantage (SEEQA, EP/Y004655/1).
@@ Appendix A Comparison to Prior Work
The qDRIFT algorithm [qDRIFT1] is a randomized Hamiltonian simulation method based on Hamiltonian sparsification. It was originally proposed for a time-independent Hamiltonian $H=\sum_{j=1}^{L}c_{j}P_{j}$ with $\|c\|_{1}=\sum_{j}|c_{j}|$. At each step, a term $P_{j}$ is sampled with probability $p_{j}=|c_{j}|/\|c\|_{1}$, and the unitary $e^{-i\,\mathrm{sgn}(c_{j})P_{j}\,\tau}$ is applied with $\tau=T\|c\|_{1}/N$, where $T$ is the total simulation time and $N$ is the number of steps. Averaging over the $N$ independent random steps yields a quantum channel whose expectation matches $e^{-iHT}$ up to an additive error $\epsilon$, with gate complexity $\mathcal{O}(\|c\|_{1}^{2}T^{2}/\epsilon)$, independent of the number of Hamiltonian terms $L$. The rotation angle $\tau$ is fixed across all steps of the algorithm but depends on the target accuracy $\epsilon$ through the choice of $N$. In contrast, the TE-PAI algorithm with constant measurement overhead achieves gate complexity $\mathcal{O}(\|c\|_{1}^{2}T^{2}/Q)$ for a tunable parameter $Q$, incurring a constant overhead factor of $\exp(Q)$. In TE-PAI, the rotation angles are fixed at $\pm\Delta$ and $\pi$, independent of the target accuracy $\epsilon$ (though dependent on $Q$). As a result, TE-PAI can achieve exact simulation with a finite gate count for any fixed $Q$. There are many variants of qDRIFT, such as for time-dependent Hamiltonians [time-dependent_l1, unified_time-dependent], which can be compared to TE-PAI in a similar manner.
@@
Next, we compare TE-PAI to a related algorithm developed in ref. [similar] which, similar to TE-PAI, simulates the time evolution by averaging over random quantum circuits. The approach decomposes the small-angle unitary rotations into linear combination of unitary matrices which are then randomly sampled to yield an unbiased estimator. In contrast, TE-PAI obtains an unbiased estimator by randomising a linear combination of superoperators in Eq. 8.